OpenStudy (anonymous):

PLEASE HELP What is the slope of the line that passes through the points (2, 8) and (6, 12)? A. –5 B. –1 C. 1 D. 5

OpenStudy (anonymous):

Well do you want just the answer or should I show the work?

OpenStudy (anonymous):

Show the work then answer

OpenStudy (black_ninja123):

Do you know the formula?

OpenStudy (anonymous):

Yea its y2-y1 over x2-x1. So Slope = 12−8 6−2 4/4 The slop is = 1

OpenStudy (black_ninja123):

Yep

OpenStudy (anonymous):

Need any more help with questions? If this is a K12 question I might have the answers.